Can anyone help me with these absolute value complex numbers problems? Thank you for your time!
Solve the following:
1) If z=5-12i then |z|=
2) If z=7-√(22i) then |z|=
√=radical or square root of 22i.

Answer by Alan3354(69443)

1) If z=5-12i
|z| = sqrt(5^2 + 12^2^) = 13
------------
2) If z=7-√(22i)
= sqrt(7^2 + 22)
= sqrt(71)